OpenStudy (anonymous):

Use a calculator to solve the equation on the interval [0,2pi]. sin3x=sinx I'm not use how to do this. help!

jimthompson5910 (jim_thompson5910):

the equation is this right? \[\Large \sin(3x) = \sin(x)\]

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

I just need to know how to do this problem on a calculator.

jimthompson5910 (jim_thompson5910):

you have something like a TI 83 right?

OpenStudy (anonymous):

IT-84

jimthompson5910 (jim_thompson5910):

TI 84 you mean?

OpenStudy (anonymous):

yes sorry.

jimthompson5910 (jim_thompson5910):

that's fine

jimthompson5910 (jim_thompson5910):

so the idea is that you can get everything to one side to get `sin(3x) - sin(x) = 0`

jimthompson5910 (jim_thompson5910):

then you type `sin(3x) - sin(x)` into y1 of the graphing calculator after you hit the `y=` key then graph

jimthompson5910 (jim_thompson5910):

you may have to adjust the window

OpenStudy (anonymous):

ok. how do I adjust the window?

jimthompson5910 (jim_thompson5910):

next to the `y=` key is the `window` key

jimthompson5910 (jim_thompson5910):

then put in something like xmin = -1 xmax = 7 ymin = -3 ymax = 3

OpenStudy (freckles):

I think an easy way to figure out the 0's is by making conjectures from the unit circle. like we see that sin(x)=sin(-x-pi) and that sin(x)=sin(x+2npi) So to solve something like \[\sin(p x)=\sin(x) \\ \text{ I think the following two equations can give you all the solutions } \\ px=x+2n \pi \\ px=-x-\pi+2n \pi \\ \text{ I'm going to rewrite that bottom equation a little } \\ px=-x+\pi(2n-1)\] then solve both equations for x. \[px=x+2 n \pi \\ \text{ can be solved by first subtracting } x \text{ on both sides } \\ px-x= 2 n \pi \\ x(p-1)=2 n \pi \\ x=\frac{2 n \pi}{p-1} \text{ last step I divided both dies by } (p-1) \\ \text{ the other equation } px=-x+\pi(2n-1) \text{ can be solved in a similar way }\]

OpenStudy (freckles):

where n is an integer

OpenStudy (anonymous):

ok so then my answers would be 0, 0.79, 2.36, 3.14, 3.93, and 5.50?

jimthompson5910 (jim_thompson5910):

@freckles has a good way to solve it algebraically yes you have the correct solutions @liefje8 see attached

OpenStudy (anonymous):

Awesome, thank you very much!!

jimthompson5910 (jim_thompson5910):

no problem